OpenStudy (shesolitt):

theres no graph. What is the slope-intercept form equation of the line that passes through (3, 4) and (5, 16)? y = 6x − 14

OpenStudy (shesolitt):

@Sepeario @pooja195

OpenStudy (pythagoras123):

Two points define a line. Let the 2 points be \[(x _{1}, y _{1}) and (x _{2}, y _{2}) \] General equation of a line is \[y-y _{1}=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } (x-x _{1})\]

OpenStudy (pythagoras123):

y-4=6(x-3)

OpenStudy (pythagoras123):

Expand and simplify to get y=6x-14

OpenStudy (shesolitt):

Thabk you so much @pythagoras123

OpenStudy (pythagoras123):

No problem!